OpenStudy (brookemariexx):

Write the equation of the function g(x) if g(x) = f(x – 2) +4 and f(x) = x3 + 2.

OpenStudy (mathmale):

First, a little housekeeping: x3 is ambiguous; I believe you mean "the cube of x," which should be written x^3.

OpenStudy (brookemariexx):

Oh yes, I meant that. My bad

OpenStudy (mathmale):

Write the equation of the function g(x) if g(x) = f(x – 2) +4 and f(x) = x^3 + 2. Step 1: replace "x" in f(x) = 2^3 +2 with (x-2). Example: f(x) becomes f(x-2).

OpenStudy (mathmale):

What do you get? f(x-2)= ??

OpenStudy (brookemariexx):

f(x-2) = 10?

OpenStudy (mathmale):

First, f(x) = x^3 + 2. Replace each 'x' with (x-2): f(x-2)=(x-2)^3 + 2. Simplify this if possible (it may not be possible).

OpenStudy (mathmale):

The problem defines g(x) as f(x-2). So, without doing any more, you could write g(x)=f(x-2)=(x-2)^3+2. OK with that?

OpenStudy (brookemariexx):

Yes, okay with that

OpenStudy (mathmale):

The problem defines g(x) as f(x-2) + 4. So, without doing any more, you could write g(x)=f(x-2)=(x-2)^3+2+4. OK with that?

OpenStudy (brookemariexx):

Yes

OpenStudy (brookemariexx):

So would g(x) = f(x-2) = (x-2)^3+6 be the final answer?

OpenStudy (mathmale):

Yes. Very good. Any questions about this process?

OpenStudy (brookemariexx):

No, thank you!

OpenStudy (mathmale):

Great pleasure. Take caer!